choppa2gage:

Jack is learning to save, spend, and give his money away. He has determined to give away 10% of his summer job earnings. If he gives away $256, how much did he earn this summer? *

jhonyy9:

let be x mean 100% 256 mean 10% cross multiplie and calcule x value x = 256 *100 /10 = ?

jhonyy9:

@XioGonz

CandyHeart13453:

gimme a sec

choppa2gage:

is this the asnwer x = 256 *100 /10 = ?

CandyHeart13453:

Multiply $256 by 10 to get 100%, then minus the 10% he gave away ($256).

CandyHeart13453:

not sure if its right tho

mw104969:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @choppa2gage is this the asnwer x = 256 *100 /10 = ? \(\color{#0cbb34}{\text{End of Quote}}\) no you gotta solve that

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 let be x mean 100% 256 mean 10% cross multiplie and calcule x value x = 256 *100 /10 = ? \(\color{#0cbb34}{\text{End of Quote}}\) I agree If 10%=256 the 100% multiplied 10 times is 2560